OpenStudy (anonymous):

1. Given the following polynomial, find a. the zeros and the multiplicity of each b. where the graph crosses or touches the x-axis c. number of turning points d. end behavior f(x)= (x-2)^3

OpenStudy (whpalmer4):

Can you tell me where \[(x-2)^3 = 0\]?

OpenStudy (anonymous):

What do you mean?

OpenStudy (whpalmer4):

What values of \(x\) make \((x-2)^3 = 0\)?

OpenStudy (anonymous):

What values of x? is the answer 0

OpenStudy (whpalmer4):

It may be helpful to remember that \((x-2)^3 = (x-2)(x-2)(x-2)\) \[(0-2)^3 = 0\]Really?

OpenStudy (anonymous):

I have no clue... I suck when it comes to math, its been 20 years sence Ive seen or dealt with math....

OpenStudy (anonymous):

But I am a quick learner, once you show me the formula, because I know its real simple, but knowing the formula, I can learn quick!!!

OpenStudy (whpalmer4):

what is 0-2?

OpenStudy (anonymous):

(2), so the value of x is 2

OpenStudy (whpalmer4):

0-2 = -2, though maybe you're an accounting person, writing (2) instead? yes, at \( x = 2, (x-2)^3 = 0\)

OpenStudy (anonymous):

okay taking notes now.. no I'm not an accountant :), I don't know why I did that....

OpenStudy (whpalmer4):

that's a typical financial report convention...but don't do it in math! all right. as I hinted, our equation can be written as \[f(x) = (x-2)^3 = (x-2)(x-2)(x-2) \] and as you found, that equals 0 at \(x = 2\). So, our zero is at \(x = 2\). Do you know or have a guess at what multiplicity means?

OpenStudy (anonymous):

multiplicity of a member of a multiset is the number of times it appears in the multiset.

OpenStudy (anonymous):

a large number or variety

OpenStudy (whpalmer4):

okay. in this context, what do you think it means?

OpenStudy (anonymous):

choice the largest number which is 2 (maybe)

OpenStudy (anonymous):

hello you still there?

OpenStudy (whpalmer4):

yes, I am. in this context, multiplicity means the number of times a root (or zero, as they are interchangeably called) is repeated. what is the multiplicity of x=2 in this equation?

OpenStudy (anonymous):

3

OpenStudy (whpalmer4):

very good. okay, now we need to figure out where the graph crosses or touches the x-axis.

OpenStudy (anonymous):

okay.....

OpenStudy (whpalmer4):

some background: if the root has multiplicity 1, that's a spot where the graph crosses the x-axis. if the root has an odd multiplicity, that's a spot where the graph crosses the x-axis. if the root has an even multiplicity, that's a spot where it touches and moves away, like in my drawing: (that's only supposed to be touching at a point, not a broad area)

OpenStudy (whpalmer4):

for our function, we have an odd multiplicity. what does that mean?

OpenStudy (anonymous):

1, or 3

OpenStudy (whpalmer4):

no, what does it mean about the behavior of our graph

OpenStudy (whpalmer4):

to be more precise, we have a zero with odd multiplicity...

OpenStudy (whpalmer4):

that zero is at x = 2 what does the graph do at x=2?

OpenStudy (anonymous):

go up 1 unit

OpenStudy (anonymous):

Im guessing I don't know........

OpenStudy (whpalmer4):

sigh. go back and read what I told you about multiplicity of roots

OpenStudy (anonymous):

okay 3

OpenStudy (whpalmer4):

I repeat: some background: if the root has multiplicity 1, that's a spot where the graph crosses the x-axis. if the root has an odd multiplicity, that's a spot where the graph crosses the x-axis. if the root has an even multiplicity, that's a spot where it touches and moves away

OpenStudy (whpalmer4):

from the multiplicity of a zero, you can determine the behavior of the graph at that value of x. at the value x = 2, where we have a zero of multiplicity 3, what will the graph do?

OpenStudy (anonymous):

gose up 3 times

OpenStudy (whpalmer4):

your only choices are "crosses the x-axis" and "touches the x-axis and retreats"

OpenStudy (anonymous):

okay whpalmer, I do thank you for your help.. but I'm really a visual learner, If your trying to teach me i need someone in person that can show me.. I'm too confused, and frustrated.. Once again thanks for your help!!!!

OpenStudy (whpalmer4):

Okay, do you know what a parabola looks like?

OpenStudy (anonymous):

like the Mcdonalds arch

OpenStudy (campbell_st):